A line has a slope of 2 and a y-intercept of (0, -4). What is the value of y
when x = 6

Answer by stanbon(75887)

A line has a slope of 2 and a y-intercept of (0, -4). What is the value of y
when x = 6
----
Equation: y = 2x - 4
---
If x = 6 , y = 2*6-4 = 12-4 = 8
============
Cheers,
Stan H.